Solve the problem. The function Y(x) = 48.94 ln can be used to estimate the number of years Y(x) after 1980 required for a certain country's population to reach x million people. In what year will the country's population reach 11 million?
2. Given that log 2 ≈ 0.301 and log 3 ≈ 0.477, find the following. log6 27

Answer by edjones(8007)

2) Given that log 2 ≈ 0.301 and log 3 ≈ 0.477, find the following. log6 27
.477 * 3 = 1.431 (log[10](27))
.301 + .477 = .778 (log[10](6))
.
Change of base formula: 1.431/.778=1.839 (log[6](27))
.
Ed